# Measuring Fitness and Importance of Species in Food Webs

### v1

@@ III Discussion
Ecosystems face intensifying threats, making it urgent to identify vulnerable species, to prevent their extinction. Current network-based approaches often reduce the complex predator-prey interactions to a single metric. While this allows us to understand which species are crucial for the functioning of a food web, it tells little to nothing about the vulnerability of species. Our method, based on a bi-dimensional non-linear map, addresses this gap, associating to each species both a importance and a fitness. The former is linked to the susceptibility to predation and measures how crucial a species is in the food web. The latter measures predatory prowess and quantifies the robustness of species. High-importance species, which often anchor lower trophic levels, appear critical to maintaining overall network stability, and their loss disproportionately impacts a wide range of other species. On the other hand, high-fitness species thrive on diverse prey pools, making them more resistant to ecosystem perturbations. Our analysis confirms this intuition, showing that importance effectively identifies species whose removal triggers significant co-extinctions. At the same time, fitness highlights species capable of persisting until the late stages of collapse, with the most vulnerable, low-fitness, species being the first to disappear. While eigenvector-in centrality can resemble our importance ranking, neither eigenvector-in nor eigenvector-out centralities adequately capture the predatory strength and resilience that define a species’ fitness. This limitation arises from the linear nature of the eigenvector centrality, which is often incapable of capturing complex, non-linear interactions.
@@
By mapping species onto a fitness-importance plane, we can reveal patterns that align with known functional roles (e.g., apex predators, base trophic nodes). While standard approaches rely on classifications by experts, our methodology is entirely algorithmic and data-driven. This framework holds clear promise for conservation and ecosystem management, where limited resources demand strategic prioritization. By jointly quantifying species’ importance and fitness, practitioners can pinpoint “keystone” nodes —those whose loss risks triggering wider co-extinctions— thereby directing scarce funding and protective measures where they have the greatest impact. Particular attention can also be devoted to those species that despite a marginal role as carbon suppliers, are in severe danger due to their low fitness. Because our algorithmic approach draws solely on quantitative interaction data, it avoids reliance on expert-based classifications and readily scales to large, complex food webs. When combined with objective risk assessment, this scalability ensures broader applicability across diverse ecosystems, ultimately supporting more targeted and cost-effective biodiversity preservation efforts.
@@ IV.1 Data
We analyzed the following seven food webs:
@@
Halodule wrightii
(shoal grass) community in Goose Creek Bay, St. Marks National Wildlife Refuge, Florida [37].
@@ IV.2 Networks Construction and Preprocessing
Food webs are represented as directed monopartite networks, where edges denote energy flow between species. In our binary model, an edge from node $i$ to node $j$ indicates that species $i$ preys upon species $j$. Self-loops are included to account for cannibalism within species. We selected six food webs for analysis and modify them to ensure irreducibility and primitivity, following the method described by [28, 40, 41]. This modification process involved:
@@
To maintain consistency across datasets, we remove three nodes (Input, Output, and Respiration) from the Florida Bay, Cypress Dry Season, and Mangrove Wet Season networks. This preprocessing step ensures that all networks represent comparable ecological entities. The macroscopic properties of these modified networks, including the number of nodes, number of links, and network density, are detailed in the Supplementary Information.
IV.3 Fitness-Importance Plane
@@
To enable cross-network comparisons, we normalize the fitness and importance values for each food web using the following procedure:
@@
    Compute fitness and importance for each species.
@@ IV.4 Extinction Areas
We implement the following species removal procedure to assess ecosystem robustness. Species are progressively removed based on rankings from specified algorithms, maintaining fixed rankings throughout the process. Cascading extinctions are monitored, with a species considered extinct when left without outgoing links or with only a self-loop. The procedure continues until all species go extinct. An extinction curve is generated by plotting the fraction of extinct species against the number of removed species. The extinction area, calculated as the integral of this curve, measures the ecosystem vulnerability. This area equals 1 when all species go extinct after the first removal and approaches 0.5 when no secondary extinctions occur. To address ranking degeneracies, we average results from 200 randomized tie-breakers and compare rankings based on importance, eigenvector-in, and in-degree centralities. For eigenvector-based rankings, we follow the method in [28]: at each step, we compute the dominant eigenvector of the column-normalized adjacency matrix and remove the highest-scoring node. The dominant eigenvalue is always 1, with its associated positive eigenvector summing to 1. For eigenvector-out rankings, we use the transpose of the adjacency matrix.
@@
Appendix S3 Fitness and Importance
S3.1 Quadrants Fitness-Importance Plane
Species in each food web are categorized into four quadrants based on their fitness and importance: Base Trophic Nodes, Sheltered Trophic Nodes, Apex Trophic Nodes, and Carbon Bridge Nodes. The full species lists for each category are provided in Appendix S6.
@@
Figure 4 reveals a strong correlation between importance and eigenvector-in centrality (Spearman coefficient = 0.93), explaining their similar performance in extinction area analysis. Conversely, Figure 5 shows a weak correlation between fitness and eigenvector-out centrality (Spearman coefficient = 0.22), accounting for significant differences in identifying resilient species.
Figure 4: Relationship between eigenvector-in centrality and importance in the Cypress food web. The Spearman correlation coefficient (s) is reported.
Figure 5: Relationship between eigenvector-out centrality and fitness in the Cypress food web. The Spearman correlation coefficient (s) is reported.
@@
Table 3 compares mean extinction areas using different ranking strategies, recomputing rankings after each species removal. The results reveal that in this case eigenvector-in centrality generally outperforms importance, which in turn outperforms in-degree. Notably, for the Coachella food web, importance shows superior performance compared to eigenvector-in centrality.
Food Web In-degree (%) Eigenvector-in (%) Importance (%)
@@
Table 3: Comparison of mean extinction area across ecosystems using different species ranking methods. We compute the in-degree, eigenvector-in, and importance rankings at the outset and recompute them after each step of the simulated extinction process. Values represent means and standard deviation. For eigenvector-in and importance, the standard deviation is smaller than the last significative digit displayed (except for the eigenvector-in in the Coachella food web, for which it is 1.0). Bold values indicate the best-performing method for each ecosystem.
Food Web In-degree (%) Eigenvector-in (%) Importance (%)
@@
Table 4: Comparison of maximum mean extinction areas across ecosystems. For each food web and ranking method (in-degree, eigenvector-in, and importance), the table presents the higher mean extinction area obtained from either the fixed or recomputed ranking strategy.
@@
Table 4 shows the maximum mean extinction areas for each food web and ranking method (in-degree, eigenvector-in, importance). The reported value represents the better result between the fixed and recomputed ranking strategies.
In Figure 6, we compare the extinction curves for all food webs under two removal strategies (recomputed and fixed initial ranking) for importance ranking. Coachella is the only food web where recomputing outperforms the fixed rankings.
Figure 6: Comparative extinction curves across food webs. Extinction trajectories for all six food webs, illustrating species loss under two removal strategies: dynamically recomputed importance ranking and fixed initial importance ranking. Each subplot depicts the proportion of extinct species versus the proportion of species removed.
Appendix S5 Robustness of Species
Table 5 presents the Spearman rank correlation coefficients between removal steps and two ecosystem metrics: average log fitness and average log eigenvector-out centrality of extinct species. These coefficients were calculated based on species removals following the importance ranking. The results demonstrate that fitness consistently outperforms eigenvector-out centrality in identifying resilient species across all food webs.
Food Web Fitness (%) Eigenvector-out (%)
@@
Table 5: Spearman rank correlation coefficients (%) between removal steps and two ecosystem metrics: average log fitness and average log eigenvector-out of extinct species. Removals were performed following the importance ranking.
@@
This superior performance of fitness holds true even when species removals are based on eigenvector-in centrality, as illustrated in Figure 7a for the Cypress food web. Figure 7b extends this analysis to all six food webs, showcasing normalized average log fitness versus normalized extinction steps. This comparison reveals consistent patterns of ecosystem resilience across diverse food web structures.
@@
Figure 7: Ecological robustness: Species removal patterns in diverse food webs. Left: Cypress food web: Average log fitness of extinct species vs. extinction step, based on eigenvector-in ranking. Labels indicate representative species becoming extinct at each step. Inset: Average log eigenvector-out centrality of extinct species vs. extinction step. Right: Normalized average log fitness vs. normalized extinction step for multiple food webs. Values are log-transformed and scaled to [0,1] for cross-ecosystem comparison.
In the specific case of the Cypress food web, Figure 8 provides additional insights into the extinction dynamics. The visualization reveals that large predators, primarily Apex Trophic Nodes, persist until later removal stages, consistent with their high fitness scores and broad prey bases.
@@
Figure 8: Cumulative species extinctions by functional category in the Cypress food web, based on importance ranking. Stacked bar chart illustrating the progressive accumulation of extinct species at each removal step. Colors represent different functional categories. Bar heights remain constant when no new extinctions occur, reflecting the cumulative nature of the data. This visualization demonstrates the differential impact of species removals on various functional groups within the ecosystem.
Appendix S6 Fitness-Importance Species Categorization
@@
Tables 6 and 7 present the species categorization for each food web based on their position in the fitness-importance plane. The Carbon Bridge Nodes for each relevant food web are as follows:

### v2

@@ III Discussion
Ecosystems face intensifying threats, making it urgent to identify vulnerable species, to prevent their extinction. Current network-based approaches often reduce the complex predator-prey interactions to a single metric. While this allows us to understand which species are crucial for the functioning of a food web, it tells little to nothing about the vulnerability of species. Our method, based on a bi-dimensional non-linear map, addresses this gap, associating to each species both an importance index and a robustness index. The former is linked to the susceptibility to predation and measures how crucial a species is in the food web. The latter measures predatory prowess and quantifies the robustness of species. High-importance index species, which often anchor lower trophic levels, appear critical to maintaining overall network stability, and their loss disproportionately impacts a wide range of other species. On the other hand, high- robustness index species thrive on diverse prey pools, making them more resistant to ecosystem perturbations. Our analysis confirms this intuition, showing that the importance index effectively identifies species whose removal triggers significant co-extinctions. At the same time, the robustness index highlights species capable of persisting until the late stages of collapse, with the most vulnerable, low- robustness index species being the first to disappear. While eigenvector-in centrality can resemble our importance index ranking, neither eigenvector-in nor eigenvector-out centralities adequately capture the predatory strength and resilience that define a species’ robustness index. This limitation arises from the linear nature of the eigenvector centrality, which is often incapable of capturing complex, non-linear interactions.
@@
By mapping species onto a robustness index-importance index plane, we can reveal patterns that align with known functional roles (e.g., apex predators, base trophic nodes). Although expert knowledge remains crucial for fully characterizing species’ functional roles, our algorithmic and data-driven approach provides a valuable, complementary layer of insight. This framework holds clear promise for conservation and ecosystem management, where limited resources demand strategic prioritization. By jointly quantifying species’ importance index and robustness index, practitioners can pinpoint “keystone” nodes —those whose loss risks triggering wider co-extinctions— thereby directing scarce funding and protective measures where they have the greatest impact. Particular attention can also be devoted to those species that despite a marginal role as carbon suppliers, are in severe danger due to their low robustness index. Because our algorithmic approach draws solely on quantitative interaction data, it readily scales to large, complex food webs. When combined with objective risk assessment, this scalability ensures broader applicability across diverse ecosystems, ultimately supporting more targeted and cost-effective biodiversity preservation efforts. However, we acknowledge that the implementation of actual conservation strategies requires the consideration of numerous additional ecological, social, and economic factors beyond the scope of this network-based analysis.
While our framework provides valuable insights into species’ roles in food webs, certain aspects require further exploration. First, basal species are represented as individual nodes but are linked to a common root node that aggregates their shared reliance on environmental resources like light, water, and nutrients. This approach simplifies the analysis by not differentiating among their specific resource dependencies and instead focuses on their collective role as energy providers. Consequently, our analysis primarily examines the vulnerability and resilience of all other species in the food web. For basal species, however, additional factors such as Grime’s classification [46]—based on competition, stress tolerance, and ruderality—may be necessary to better capture their ecological resilience and responses to cascading effects. Second, species such as lizards in the Cypress food web appear peripheral in our analysis due to their limited predatory skills and resistance to predation. This could reflect biological realities, such as population constraints driven by nesting site availability rather than trophic interactions. However, it may also result from aggregation biases in network construction, where diverse prey groups are aggregated into single nodes, masking their ecological importance. To preliminarily assess the influence of network resolution, we conducted node aggregation experiments. Our initial findings indicate that the impact of such aggregation on our metrics of ecosystem fragility is not uniform. Aggregating multiple basal trophic nodes yielded the most pronounced effects, suggesting that the resolution at lower trophic levels could be particularly critical for accurately assessing species’ roles using our framework. Further details on these aggregation experiments and their results can be found in Section 6 of the Supplementary Information. Systematic investigation into the sensitivity of our results to food web resolution, across different aggregation strategies and trophic levels, represents a valuable direction for future research. Third, our analysis relies on a static and binary representation of food webs, which inherently simplifies the dynamic and often weighted nature of real ecological interactions. The species removal simulations, while useful for comparative purposes, also present a somewhat idealized view of extinction cascades, as they do not incorporate adaptive behaviors like prey switching or the continuous changes in interaction strengths that characterize natural ecosystems. Finally, while our initial study focused on demonstrating the methodology across a diverse set of six existing food webs, future research could readily expand this analysis to the larger collection of currently available food web datasets to assess the generality of our findings. Subsequently, experimental validation remains a critical avenue, where collecting data on real-world ecological networks under anthropogenic disturbances could provide empirical evidence for the predictive power of the proposed framework and help quantify species’ actual endangerment levels across their ecosystems.
@@ IV.1 Data
We analyzed the following six food webs:
@@ IV.2 Networks Construction and Preprocessing
Food webs are represented as directed monopartite networks, where edges denote energy flow between species. In our binary model, an edge from node $i$ to node $j$ indicates $i$ preys upon species $j$, with no weights assigned to the links. Self-loops are included to account for cannibalism within species. We selected six food webs for analysis and modified them to ensure irreducibility and primitivity, following the method described by [29, 50, 51]. This modification process involved:
@@
We remove three nodes (Input, Output, and Respiration) from the Florida Bay, Cypress Dry Season, and Mangrove Wet Season networks to maintain consistency across datasets. For the St. Marks food web, only step 3 (add the links between root and all other nodes) is performed, since the first two were already present in the data. Furthermore, we manually added one missing link, comparing the data with the adjacency matrix present in the original paper [47]. These preprocessing steps ensure that all networks represent comparable ecological entities. The macroscopic properties of these modified networks, including the number of nodes, links, and network density, are detailed in the Supplementary Information.
IV.3 Robustness index-Importance index Plane
@@
To enable cross-network comparisons, we normalize the robustness index and importance index values for each food web using the following procedure:
@@
    Compute the robustness index and importance index for each species.
@@ IV.4 Extinction Areas
We implement the following species removal procedure to assess ecosystem robustness. Species are progressively removed based on rankings from specified algorithms, maintaining fixed rankings throughout the process. Cascading extinctions are monitored, with a species considered extinct when left without outgoing links or with only a self-loop. The procedure continues until all species go extinct. An extinction curve is generated by plotting the fraction of extinct species against the number of removed species. The extinction area, calculated as the integral of this curve, measures the ecosystem vulnerability. This area equals 1 when all species go extinct after the first removal and approaches 0.5 when no secondary extinctions occur. When ranking species for removal, multiple species can sometimes receive the same rank value (a situation known as rank degeneracy). To address this, we employ a randomized tie-breaker. When ties occur, we randomly select which of the equally ranked species to remove. To account for the variability introduced by these random choices, we average the results from 200 repetitions of the entire removal process. In each repetition, ties are broken randomly. We use this approach to compare rankings based on our importance index, eigenvector in-centrality, and in-degree centrality. For eigenvector-based rankings, we follow the method in [29]: at each step, we compute the dominant eigenvector of the column-normalized adjacency matrix and remove the highest-scoring node. The dominant eigenvalue is always 1, with its associated positive eigenvector summing to 1. For eigenvector-out rankings, we use the transpose of the adjacency matrix.
Author Contributions
Conceptualization: G.D.M. and V.D.P.S.; methodology: E.C., G.D.M. and V.D.P.S.; software: E.C.; validation: E.C., G.D.M. and V.D.P.S.; formal analysis: E.C. and G.D.M.; investigation: E.C. and G.D.M.; data curation: E.C.; writing—original draft preparation: E.C. and G.D.M.; writing—review and editing: E.C., G.D.M. and V.D.P.S.; visualization: E.C. and G.D.M.; supervision: G.D.M. and V.D.P.S.; project administration: V.D.P.S. All authors have read and agreed to the published version of the manuscript.
Competing Interests
The authors declare no competing interests.
Data Availability
The edgelists and species’ names for St.Marks, Coachella, and Benguela food webs were constructed by examining the adjacency matrices in their original publications [47, 48, 49]. The edgelists and species’ names for the Florida Bay, Cypress Dry Season, Mangrove Wet Season food webs [45] were originally available at [52], sourced from [53].
All the edgelists and species’ names are available in the Zenodo repository [54].
Acknowledgments
This work was funded by the City of Vienna, Municipal Department 7, and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Energy, Mobility, Innovation and Technology as part of the project GZ 2021-0.664.668.
We thank Fariba Karimi, Rafael Prieto-Curiel, and Andrea Perna.
Declaration of Generative AI and AI-assisted technologies in the writing process
We acknowledge the use of LLMs for text refining. The authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@ Acknowledgments
Appendix S3 Robustness index and Importance index
S3.1 Quadrants Robustness index - Importance index Plane
Species in each food web are categorized into four quadrants based on their robustness index and importance index: Base Trophic Nodes, Sheltered Trophic Nodes, Apex Trophic Nodes, and Carbon Bridge Nodes. The full species lists for each category are provided in Appendix S7.
@@
Figure 4 reveals a strong correlation between the importance index and eigenvector-in centrality (Spearman coefficient = 0.93), explaining their similar performance in extinction area analysis. Conversely, Figure 5 shows a weak correlation between the robustness index and eigenvector-out centrality (Spearman coefficient = 0.22), accounting for significant differences in identifying resilient species.
Figure 4: Relationship between eigenvector-in centrality and importance index in the Cypress food web. The Spearman correlation coefficient (s) is reported.
Figure 5: Relationship between eigenvector-out centrality and robustness index in the Cypress food web. The Spearman correlation coefficient (s) is reported.
@@
Table 3 compares mean extinction areas using different ranking strategies, recomputing rankings after each species removal. The results reveal that in this case eigenvector-in centrality generally outperforms the importance index, which in turn outperforms in-degree. Notably, for the Coachella food web, the importance index shows superior performance compared to eigenvector-in centrality.
Food Web In-degree (%) Eigenvector-in (%) Importance index (%)
@@
Table 3: Comparison of mean extinction area across ecosystems using different species ranking methods. We compute the in-degree, eigenvector-in, and importance index rankings at the outset and recompute them after each step of the simulated extinction process. Values represent means and standard deviation. The standard deviation for the eigenvector-in and importance index is smaller than the last significative digit displayed (except for the eigenvector-in in the Coachella food web, for which it is 1.0). Higher percentages indicate greater ecosystem sensitivity to species loss based on the respective ranking method. Bold values indicate the best-performing method for each ecosystem.
Food Web In-degree (%) Eigenvector-in (%) Importance index (%)
@@
Table 4: Comparison of maximum mean extinction areas across ecosystems. For each food web and ranking method (in-degree, eigenvector-in, and importance index), the table presents the higher mean extinction area obtained from either the fixed or recomputed ranking strategy.
@@
Table 4 shows the maximum mean extinction areas for each food web and ranking method (in-degree, eigenvector-in, importance index). The reported value represents the better result between the fixed and recomputed ranking strategies.
In Figure 6, we compare the extinction curves for all food webs under two removal strategies (recomputed and fixed initial ranking) for importance index ranking. Coachella is the only food web where recomputing outperforms the fixed ranking.
Figure 6: Comparative extinction curves across food webs. Extinction trajectories for all six food webs, illustrating species loss under two removal strategies: dynamically recomputed importance index ranking and fixed initial importance index ranking. Each subplot depicts the proportion of extinct species versus the proportion of species removed, highlighting differences between adaptive and static removal processes.
Food Web Robustness index (%) Eigenvector-out (%)
@@
Table 5: Spearman rank correlation coefficients (%) between removal steps and ecosystem metrics: average log robustness index and average log eigenvector-out of extinct species. Removals were performed following the importance index ranking.
Appendix S5 Robustness of Species
Table 5 presents the Spearman rank correlation coefficients between removal steps and two ecosystem metrics: average log robustness index and average log eigenvector-out centrality of extinct species. These coefficients were calculated based on species removals following the importance index ranking. The results demonstrate that the robustness index consistently outperforms eigenvector-out centrality in identifying resilient species across all food webs.
@@
This superior performance of the robustness index holds true even when species removals are based on eigenvector-in centrality, as illustrated in Figure 7a for the Cypress food web. Figure 7b extends this analysis to all six food webs, showcasing normalized average log robustness index versus normalized extinction steps. This comparison reveals consistent patterns of ecosystem resilience across diverse food web structures.
@@
Figure 7: Ecological robustness: Species removal patterns in diverse food webs. Left: Cypress food web: Average log robustness index of extinct species vs. extinction step, based on eigenvector-in ranking. Labels indicate representative species becoming extinct at each step. Inset: Average log eigenvector-out centrality of extinct species vs. extinction step. Right: Normalized average log robustness index vs. normalized extinction step for multiple food webs. Values are log-transformed and scaled to [0,1] for cross-ecosystem comparison. This plot illustrates ecosystem resilience patterns and differential impacts of species loss across varied food web structures.
@@
Figure 8: Cumulative species extinctions by functional category in the Cypress food web, based on the importance index ranking. Stacked bar chart illustrating the progressive accumulation of extinct species at each removal step. Colors represent different functional categories. Bar heights remain constant when no new extinctions occur, reflecting the cumulative nature of the data. This visualization demonstrates the differential impact of species removals on various functional groups within the ecosystem.
In the specific case of the Cypress food web, Figure 8 provides additional insights into the extinction dynamics. The visualization reveals that large predators, primarily Apex Trophic Nodes, persist until later removal stages, consistent with their high robustness index scores and broad prey bases.
Appendix S6 Aggregation Analysis
In the Florida Bay food web, we aggregated eight phytoplankton species ($2\mu$m Spherical Phytoplankt, Synedococcus, Oscillatoria, Small Diatoms ($<20\mu$m), Big Diatoms ($>20\mu$m), Dinoflagellates, Other Phytoplankton, Benthic Phytoplankton) into a single “Phytoplankton” node. We also aggregated two heron categories (Big Herons & Egrets, Small Herons & Egrets) into a single “Herons & Egrets” node, three turtle species (Loggerhead Turtle, Green Turtle, Hawksbill Turtle) into “Turtles”, and three killifish species (Other Killifish, Goldspotted killifish, Rainwater killifish) into “Killifish”. We then examined the effect of these individual aggregations on the total extinction area resulting from simulated species removals and cascading effects. Aggregating the eight phytoplankton species led to the most substantial increase in the extinction area (eigenvector centrality-based area to 91.0, importance-based area to 90.1). The aggregation of Herons & Egrets resulted in a minor change. Aggregating turtles led to a slight decrease in the eigenvector centrality-based extinction area (to 82.1) and a minor increase in the importance-based area (to 82.4). Aggregating killifish resulted in a slight decrease in both the eigenvector centrality-based extinction area (to 82.3) and the importance-based area (to 82.1).
Appendix S7 Species Categorization
@@
Tables 6 and 7 present the species categorization for each food web based on their position in the robustness index-importance index plane. The Carbon Bridge Nodes for each relevant food web are as follows: